A certain distribution of money is distribution to three boys in the ratio of 4:10:18 the difference between largest share and smallest share is 84000. Find the money to be distribution.
Answer by ankor@dixie-net.com(22740)

A certain distribution of money is distribution to three boys in the ratio of 4:10:18
let let x = the multiplier
then to total amt given to the three boys
4x + 10x + 18x = 32x
:
the difference between largest share and smallest share is 84000.
18x - 4x = 84000
14x = 84000
x = 84000/14
x = 6000 is the multiplier
:
Find the money to be distributed.
32x = total amt
32 * 6000 = $192,000
:
:
Check this, find the actual amt each boy got:
6000(4) = 24000
6000(10) = 60000
6000(18) = 108000
---------------------
total amt: 192000